OpenStudy (sgayton27):

Question Part Points Submissions Used 1 0/1 3/100 Total 0/1 Use the unit circle to find all values of θ between 0 and 2π for the following; tanθ = (-sqrt3)/3

OpenStudy (sgayton27):

\[\tan \theta= -\frac{ \sqrt{3} }{ 3 }\]

OpenStudy (solomonzelman):

Agree!